### v2

# Optimal Qubit Purification and Unitary Schur Sampling via Random SWAP Tests

Abstract
The goal of qubit purification is to combine multiple noisy copies of an unknown pure quantum state to obtain one or more copies that are closer to the pure state. We show that a simple protocol based solely on random SWAP tests achieves the same fidelity as the Schur transform, which is optimal. This protocol relies only on elementary two-qubit SWAP tests, which project a pair of qubits onto the singlet or triplet subspaces, to identify and isolate singlet pairs, and then proceeds with the remaining qubits. For a system of $n$ qubits, we show that after approximately $T\approx n\ln n$ random SWAP tests, a sharp transition occurs: the probability of detecting any new singlet decreases exponentially with $T$. Similarly, the fidelity of each remaining qubit approaches the optimal value given by the Schur transform, up to an error that is exponentially small in $T$. More broadly, this protocol achieves what is known as weak Schur sampling and unitary Schur sampling with error $\epsilon$, after only $2n\ln(n\epsilon^{-1})$ SWAP tests. That is, it provides a lossless method for extracting any information invariant under permutations of qubits, making it a powerful subroutine for tasks such as quantum state tomography and metrology.
I Introduction
A broad and fundamental class of problems in quantum information science involves extracting and processing information encoded in the quantum state of many qubits. A central obstacle to realizing quantum advantage in these settings is the efficient implementation of the necessary information-extraction operations (e.g., using only a limited set of experimentally accessible, elementary operations).
In many of these applications, the goal is to extract information that does not depend on the order of the given qubits-that is, permutationally invariant (PI) information. This includes, for instance, quantum metrology and state tomography [helstrom1969quantum, holevo2011probabilistic, giovannetti2011advances, giovannetti2006quantum, TomographyHarrow], where the goal is to use many copies of a quantum state to estimate an unknown parameter or reconstruct a classical description of the state. Another example discussed further below is qubit purification [Cirac], where many noisy copies of an unknown pure state are combined to obtain one or more copies with higher fidelity. Other important examples of applications include density matrix exponentiation [densitylloyd], quantum emulation [marvian2016universal], coherence distillation [marvian2020coherence], and majority voting [majority].
Overview of Results
In this work, we show that a simple protocol can efficiently and optimally extract the PI information from systems composed of an arbitrarily large number of qubits. Remarkably, the protocol requires only the SWAP test, a simple and widely used two-qubit measurement that determines whether a pair of qubits is in the singlet state $(|01\rangle-|10\rangle)/\sqrt{2}$ or in the orthogonal triplet subspace. The protocol operates by performing SWAP tests on randomly chosen pairs of qubits. Whenever a pair is found to be in the singlet state, it is discarded, and the protocol proceeds with the remaining qubits (see Fig. 1). The protocol is lossless in the following sense: from its output, one can recover the original input state up to an unknown permutation of the qubits. In this way, the protocol gradually removes entropy from the system while preserving its PI information content.
We show that for a system of $n$ qubits, after approximately $2n\ln n$ SWAP tests, a sharp transition occurs: with high probability, no additional singlets are found, and the state of the remaining qubits converges to a fixed state within their totally symmetric subspace. More formally, we compare the output of this protocol with that of the so-called Schur transform [weakchilds, SchurTransformHarrow, harrowthesis], a powerful tool in quantum information with broad applications, which, in particular, has been commonly used for extracting PI information. We prove that for PI states, the trace distance between the output of the Schur transform and the protocol based on SWAP tests is bounded by $\epsilon>0$, provided that $T$, the number of SWAP tests, satisfies
$T\geq 2n\ln(n\epsilon^{-1})\ .$ (1)
For a system with $n$ qubits, clearly at least $\tfrac{n}{2}$ SWAP tests are needed to interact with all qubits at least once. Therefore, roughly speaking, our result shows that going beyond this minimum by a factor of $\ln n$ is sufficient to implement the Schur transform on PI states, with an error that is exponentially small.
Figure 1: Qubit Purification and Unitary Schur Sampling via Random SWAP Tests. This protocol uses random SWAP tests to detect and isolate pairs of qubits in the singlet state $(|01\rangle-|10\rangle)/\sqrt{2}$. In the figure above, the partitions $B_{k}$, for $k=1,\dots$, correspond to the detected and separated singlets. The partitions $A_{k}$, for $k=1,\dots$, contain the remaining qubits. At each step $k$, we randomly select a pair of qubits from partition $A_{k}$, perform a SWAP test on them, and if the outcome projects onto the singlet state, we move that pair to partition $B_{k}$ and proceed to step $k+1$. Otherwise, if the SWAP test does not detect a singlet, we randomly choose another pair from $A_{k}$ and repeat the process. We prove that after $T\geq 2n\ln(n\epsilon^{-1})$ SWAP tests, with probability at least $1-\epsilon$, all singlets in the system have been detected. Furthermore, for the initial state $\rho^{\otimes n}$ with $\rho=(1-p)|\psi\rangle\langle\psi|+p\mathbb{I}/2$, where $|\psi\rangle$ is unknown, any qubit in $A_{k}$ will have the highest achievable fidelity given in Eq. (2), up to an additional error of at most $\epsilon$.
As an example of applications, we demonstrate that this protocol for state purification [Cirac]. Namely, we apply it to $\rho^{\otimes n}$, i.e., $n$ copies of state $\rho=(1-p)|\psi\rangle\langle\psi|+p\mathbb{I}/2$, where $|\psi\rangle$ is an unknown pure state and $0<p<1$, in order to obtain a single qubit with higher fidelity with respect to $|\psi\rangle$. Cirac et al. [Cirac] have shown how the optimal fidelity can be achieved using the Schur transform, and the fidelity of the output $\rho_{\text{opt}}(n)$ with the desired output state $|\psi\rangle$ is given by
$\langle\psi|\rho_{\text{opt}}(n)|\psi\rangle=1-\frac{1}{2n}\,\frac{p}{(1-p)^{2}}+\mathcal{O}\left(\frac{1}{n^{2}}\right)\ .$ (2)
Our result then reveals that, up to an arbitrarily small additional error $\epsilon$, the same fidelity can be achieved using $2n\ln(n\epsilon^{-1})$ random SWAP tests, by simply returning one of the qubits that has not been discarded by the SWAP test. See Fig. 2 for an example, and Sec. LABEL:Sec:App:Pur for further details.
As we further explain below, the working principle behind this purification protocol is simple: for PI states, the only relevant information is encoded in the reduced state of the $\mathrm{SU}(2)$ irreducible representations (irreps). Since the SWAP test on each pair of qubits respects this symmetry, the initial state can always be recovered after any number of SWAP tests, regardless of the specific choices of measured qubits and the outcomes. By discarding the detected singlets and continuing the SWAP tests on the remaining qubits, we gradually detect all the singlets in the system. This ultimately yields a state restricted to the totally symmetric subspace of the remaining qubits, where the SWAP test no longer detects any singlets. In this case, by discarding all the remaining qubits except one, we achieve the optimal purification limit in Eq.(2).
We note that, prior to this work, Childs et al. [Childs2025streamingquantum] used sequential SWAP tests to implement qubit purification. However, in their approach, when a SWAP test identifies a pair of qubits in the triplet subspace, one qubit is retained and forwarded to the next level, and the other is discarded. This simplifies the analysis, as the remaining qubits are fully uncorrelated, but it leads to wasted resources and thus suboptimal fidelity as well as suboptimal gate complexity (see Sec.LABEL:Sec:App:Pur).
We also note that in the context of symmetry-protected quantum computation, Freedman et al. [freedman2021symmetry] showed in 2021 that the outcomes of random SWAP tests, referred to as triplet/singlet measurements in their work, determine the total angular momentum of a system of qubits. Building on this, Rudolph and Virmani [rudolph] showed in 2023 that, starting from qubits in the maximally mixed state, one can, by repeatedly measuring angular momentum in this way, gradually grow a subset of qubits whose joint state is restricted to their totally symmetric subspace, namely, in the maximally mixed state on this subspace. (At each step, one performs random SWAP tests between the previously obtained subset and a new qubit in the maximally mixed state).
Our results reveal broader applications of random SWAP tests, namely for performing the Schur transform on PI states, which is useful, for example, in the context of state purification and tomography, and they provide more precise and tighter error analysis, uncovering a sharp threshold phenomenon around $T_{\ast}\approx 2n\ln n$.  [^1] Rudolph and Virmani [rudolph] show that the convergence can be achieved to the maximally mixed state on the totally symmetric subspace in $\approx n^{2}$ steps. Freedman et al. [freedman2021symmetry] briefly discuss using random SWAP tests for measuring angular momentum $j$, as a useful subroutine in the so-called symmetry-protected quantum computation. They argue that “since spin is quantized, the convergence is actually faster than polynomial. Once the number of measurements is polynomially large compared to $n^{2}$, the convergence in accuracy becomes exponential.” We also note that, in the special case of $n=3$ qudits, it was previously observed in [li2024optimal] that unitary Schur sampling can be achieved via sequential SWAP tests.
Figure 2: The fidelity of the output qubit of this purification protocol, with the desired ideal pure state, after $T$ SWAP tests. Initially the system contains $n=100$ copies of $\rho=(1-p)|\psi\rangle\langle\psi|+p\mathbb{I}/2$. As the number of SWAP tests grows, the fidelity of the output state converges to the optimal fidelity given by Eq.(2). The vertical dotted line corresponds to $n\ln n$, which roughly indicates the order of the number of required SWAP tests. See Sec.LABEL:Sec:App:Pur for further details.
II Setup
Consider a system of $n$ qubits with the total Hilbert space $(\mathbb{C}^{2})^{\otimes n}$. We are interested in the properties of quantum states that are independent of the order of the qubits. More formally, for any state $\sigma$, we are interested in the information that can be obtained from the PI state $\mathcal{P}(\sigma)$, where $\mathcal{P}$ is the quantum channel defined by applying a permutation on $n$ qubits, chosen uniformly at random from the group of all such permutations $\mathbb{S}_{n}$ (See Appendix LABEL:App:SU(2) for further details). By applying such random permutations, any state $\sigma$ is projected to its symmetrized version $\mathcal{P}(\sigma)$. Hence, in the following, we often restrict our attention to the case of PI states.
As we discuss further below, thanks to Schur-Weyl duality, such states can be conveniently characterized using the representation theory of the group $\mathrm{SU}(2)$, i.e., single-qubit unitaries with determinant 1, which is widely used in physics (e.g., to describe rotations). Recall that under the action of $\mathrm{SU}(2)$ symmetry, $U\mapsto U^{\otimes n}$ with $U\in\mathrm{SU}(2)$, the Hilbert space decomposes into subspaces corresponding to inequivalent irreps of $\mathrm{SU}(2)$. These are commonly labeled by the eigenvalues $j(j+1)$ of the squared angular momentum operator (also known as the Casimir operator) $J^{2}=J_{x}^{2}+J_{y}^{2}+J_{z}^{2}$, where $J_{w}=\sum_{i=1}^{n}\sigma_{w}^{(i)}/2\quad\text{for }w\in\{x,y,z\}.$ The possible values of $j$ are integers $j=0,\ldots,n/2$ when $n$ is even, and half-integers $j=1/2,\ldots,n/2$ when $n$ is odd.
Then, under the action of $\mathrm{SU}(2)$ transformations, the total Hilbert space of qubits decomposes as
$(\mathbb{C}^{2})^{\otimes n}\cong\bigoplus_{j=j_{\text{min}}}^{n/2}\big(\mathbb{C}^{2j+1}\otimes\mathbb{C}^{m(n,j)}\big)\ ,$ (3)
where $\mathrm{SU}(2)$ acts irreducibly on $\mathbb{C}^{2j+1}$ with multiplicity $m(n,j)={\tbinom{n}{\frac{n}{2}-j}}\times\frac{2j+1}{\frac{n}{2}+j+1}$. According to Schur-Weyl duality, the group of unitaries generated by pairwise SWAPs of qubits, which is a representation of $\mathbb{S}_{n}$, is block-diagonal with respect to this decomposition, acts trivially on the subsystem $\mathbb{C}^{2j+1}$ and acts irreducibly on $\mathbb{C}^{m(n,j)}$. Indeed, each $\mathbb{C}^{m(n,j)}$ with a different $j$ corresponds to an inequivalent irrep of this group. In particular, the $(n+1)$-dimensional subspace corresponding to the maximum angular momentum $j=n/2$, also known as the totally symmetric subspace, carries the trivial representation of the permutation group $\mathbb{S}_{n}$. This subspace includes all pure states that are invariant under permutations, such as GHZ and Dicke states.
It follows that any PI $n$-qubit state $\sigma$, such as $\sigma=\rho^{\otimes n}$, is block-diagonal with respect to this decomposition and takes the form
$\sigma=\bigoplus_{j=j_{\text{min}}}^{n/2}p_{j}\left(\sigma_{j}\otimes\frac{\mathbb{I}_{m(n,j)}}{m(n,j)}\right)\ ,$ (4)
where $p_{j}=\operatorname{Tr}(\sigma\Pi_{j})$ is the probability that the state $\sigma$ is found in the subspace with total angular momentum $j$, i.e., the subspace $\mathbb{C}^{2j+1}\otimes\mathbb{C}^{m(n,j)}$, $\Pi_{j}$ denotes the projector onto this subspace, and $\sigma_{j}$ is a density operator on $\mathbb{C}^{2j+1}$, the irrep of $\mathrm{SU}(2)$ with angular momentum $j$.
III Unitary Schur Sampling
For many applications in quantum information and computation, it is useful to physically realize the transformation that, for a given PI input state $\sigma$, measures $j$, i.e., returns the angular momentum value $j$ with probability $p_{j}$, along with the corresponding post-measurement state $\sigma_{j}$. In [cervero2024memory], this operation is called unitary Schur sampling.
In practice, for many applications, such as purification [Cirac] and coherence distillation [marvian2020coherence], rather than returning the post-measurement state $\sigma_{j}$ as a subsystem with dimension $2j+1$, it is often more convenient and useful to have an output system whose Hilbert space is of constant size. A natural and convenient way to achieve this is to encode the state $\sigma_{j}$ in the $(2j+1)$-dimensional totally symmetric subspace of $2j$ qubits. The remaining n2jn-2j qubits (which is always an even number) can be prepared in pure singlet states, i.e., in the state $|\xi\rangle=(|01\rangle-|10\rangle)/\sqrt{2}$ (As we see below, this choice is made to respect the $\mathrm{SU}(2)$ symmetry). That is, the goal is to realize the transformation
$\mathcal{E}_{\text{Schur}}(\sigma)=\sum_{j=j_{\text{min}}}^{n/2}\widetilde{\sigma}_{j}\otimes\xi^{\otimes(\tfrac{n}{2}-j)}\otimes|j\rangle\langle j|_{\text{reg}}\ ,$ (5)
where $\xi=|\xi\rangle\langle\xi|$, $\widetilde{\sigma}_{j}$ is an unnormalized density operator in the totally symmetric subspace of $2j$ qubits satisfying $\operatorname{Tr}(\widetilde{\sigma}_{j})=p_{j}$, and we have appended the system with an additional (classical) register that keeps track of the value of $j$ as $|j\rangle_{\text{reg}}$. [^2] Obviously, this register does not need to be a separate quantum device; it could be, for example, the controller of the experimental apparatus or simply the experimentalist’s memory! As we further explain below in Sec.LABEL:Sec:compress, if one is interested in obtaining state $\sigma_{j}$ encoded in $\lceil\log_{2}(2j+1)\rceil$ qubits, this can be achieved using an extra compression circuit, which converts $\widetilde{\sigma_{j}}/p_{j}$ to $\sigma_{j}$.
To explicitly write the matrix elements of $\widetilde{\sigma}_{j}$, we first consider a basis for the totally symmetric subspace of $2j$ qubits, namely, the states $|j,m\rangle$ for $m=-j,\dots,j$, the normalized eigenvectors of $\sum_{i=1}^{2j}\sigma_{z}^{(i)}/2$ in this subspace, also known as the Dicke states. In particular, $|j,m\rangle$ is the uniform superposition of all computational basis states in $\{|0\rangle,|1\rangle\}^{\otimes 2j}$ that contain $j+m$ qubits in state $|0\rangle$ and jmj-m qubits in state $|1\rangle$ (see Appendix LABEL:App:SU(2) for the explicit formula). Then,
$|j,m\rangle\otimes|\xi\rangle^{\otimes(\tfrac{n}{2}-j)}\ \ ,\ \ \ \ \ \text{for}\ \ \ m=-j,\cdots,j\ ,$
is a basis for a single copy of an irrep of $\mathrm{SU}(2)$ with angular momentum $j$ in $\mathbb{C}^{2j+1}\otimes\mathbb{C}^{m(n,j)}$. More precisely, with respect to this decomposition, these vectors span the subspace $\mathbb{C}^{2j+1}\otimes|\eta_{0}\rangle$, where $|\eta_{0}\rangle$ is a fixed state in $\mathbb{C}^{m(n,j)}$, the multiplicity subsystem associated to irrep $j$ (See Appendix LABEL:App:SU(2) for further discussion).
Then, according to Eq.(4), the non-zero matrix elements of $\widetilde{\sigma}_{j}$ are determined by the matrix element of $\sigma$ via [^3] It is worth noting that for state $\sigma=\rho^{\otimes n}$, the corresponding $\widetilde{\sigma}_{j}$, up to a normalization, is independent of $n$, i.e., $\frac{\langle j,m^{\prime}|\widetilde{\sigma}_{j}|j,m\rangle}{m(n,j)}=\left[\frac{1-\operatorname{Tr}(\rho^{2})}{2}\right]^{\tfrac{n}{2}-j}\times\langle j,m^{\prime}|\rho^{\otimes 2j}|j,m\rangle\ .$ (6) Therefore, while in general, the probability of observing angular momentum $j$ is determined by the number of given copies of state $\rho$ and its purity $\operatorname{Tr}(\rho^{2})$, the corresponding reduced state in the sector with angular momentum $j$ is independent of $n$.
$\frac{\langle j,m^{\prime}|\widetilde{\sigma}_{j}|j,m\rangle}{m(n,j)}=\big[\langle j,m^{\prime}|\langle\xi|^{\otimes(\tfrac{n}{2}-j)}\big]\sigma\big[|j,m\rangle|\xi\rangle^{\otimes(\tfrac{n}{2}-j)}\big]\ .$ (7)
Eq. (5), together with Eq. (7), fully defines the action of the channel $\mathcal{E}_{\text{Schur}}$ on any PI state. Explicitly,
$\mathcal{E}_{\text{Schur}}(\sigma)=\sum_{j=j_{\text{min}}}^{n/2}M_{j}\sigma M_{j}\otimes|j\rangle\langle j|_{\text{reg}}\ ,$ (8)
where
$M_{j}=\sqrt{m(n,j)}\left(\sum^{j}_{m=-j}|j,m\rangle\langle j,m|\right)\otimes\xi^{\otimes(\tfrac{n}{2}-j)}\ .$ (9)
This definition can be extended to a general $n$-qubit state $\sigma$ by first applying the channel $\mathcal{P}$, which projects any input state to a PI state, and then applying the channel in Eq. (5) to the resulting symmetrized state.